OpenStudy (anonymous):

A sequence is defined by the recursive formula f (n + 1) = f(n) – 2. If f(1) = 18, what is f(5)?

OpenStudy (anonymous):

Start off by plugging in 1 into recursion. What would that give you?

OpenStudy (anonymous):

i would say 80

OpenStudy (anonymous):

Well, if you plug in n = 1, youd have this: \(f(1+1) = f(1) - 2\) That's what I meant to do

OpenStudy (anonymous):

So just to finish off the idea: \(f(1+1) = f(1) - 2\) \(f(2) = f(1) - 2\) But f(1) is 18 as the problem told us, so \(f(2) = 18 -2\) \(f(2) = 16\) That make sense what I did?

OpenStudy (anonymous):

omg thank you

OpenStudy (anonymous):

i hate math tbh

OpenStudy (anonymous):

Most people do, lol. But yeah, the next thing youd do is plug in 2: \(f(2+1) = f(2) - 2\) \(f(3) = f(2) - 2\) \(f(3) = 16-2\) \(f(3) = 14\) So the idea is Im going to keep doing this until I get to f(5). Or you might be able to see the pattern and predict what f(5) would be from here. Either way, think you got it from here?